OpenStudy (anonymous):

The formula for determining interest compounded monthly is A = P(1 + r over 12)12t, where A represents the amount invested after t years, P the principal invested, and r the interest rate. Jimmy invests $1,500 at an interest rate of 10% for 4 years, while Jenny invests $1,500 at an interest rate of 5% for 6 years. Determine the amount of return gained by Jimmy and Jenny. Can someone check my work?

OpenStudy (mathstudent55):

Where's your work? I only see the question.

OpenStudy (anonymous):

can i get the explanation ...formula is....r=(1+i/n)^n-1...

OpenStudy (whpalmer4):

\[A = P(1+\frac{r}{N})^{Nt}\]Here A is the future amount, P is the principal or present value, r is the annual interest rate expressed as a decimal, N is the number of compounding periods per year (here 12), and t is the number of years. After one compounding period, the total is \(P * (1+r/N= P + Pr/N\). After two compounding periods, the total is \(P(1+r/N)*(1+r/N)\)\) and so on.

OpenStudy (whpalmer4):

So you'll do the computation twice: once with P = 1500, r = 0.1, N = 12, t = 4 (the Jimmy case) and once with P = 1500, r = 0.05, N = 12, t = 6 (the Jenny case) Post what you got for the two and I'll tell you if I agree.

OpenStudy (anonymous):

ya...i know all p,r n and t.....jst need to know how to solve it....A=1500(1+0.1/12)^12*4

OpenStudy (whpalmer4):

So just put the numbers in and grind it out...

OpenStudy (whpalmer4):

what is (1+0.1/12)? What is that raised to the (12*4) power? What do you get when you multiply that by 1500? (btw, ^12*4 is not correct — exponentiation has higher precedence than multiplication, so writing x^12*4 = (x^12)*4, not x^48)

OpenStudy (anonymous):

k...thanxs

OpenStudy (anonymous):

(1+1/120)=121/120....is this right....? for (1+0.1/12)

OpenStudy (whpalmer4):

Yes, you could do it that way.

OpenStudy (whpalmer4):

12.1/12 works as well.

OpenStudy (anonymous):

ya....so how to calculate the power i.e...(12*4)...to this

OpenStudy (anonymous):

121/120

OpenStudy (whpalmer4):

well, do you have a calculator?

OpenStudy (anonymous):

ya...

OpenStudy (whpalmer4):

4 function, or scientific or financial?

OpenStudy (whpalmer4):

if anything but a 4 function calc, there's probably a y^x button or something equivalent

OpenStudy (anonymous):

scientific

OpenStudy (whpalmer4):

okay, there will definitely be a button that does exponentiation on there.

OpenStudy (anonymous):

k..without calcy we can't do it.....